# Rigorous Test for Quantum Integrability and Nonintegrability

### v2

@@ II.2 Main Result
We have the following isomorphisms between linear spaces:
$\mathcal{B}_{\leq}^{(2)}$ $\cong\mathcal{B}_{\leq}^{(3)}\cong\dots\cong\mathcal{B}_{\leq}^{(N/2)},$ (7)
$\mathcal{B}_{<}^{(3)}$ $\cong\mathcal{B}_{<}^{(4)}\cong\dots\cong\mathcal{B}_{<}^{(N/2)},$ (8)
where the actions of the isomorphisms in Eq. (8) on $\mathcal{B}_{\leq}^{(k)}$ are obtained by the restriction of the domain of the isomorphisms (7).
These isomorphisms yield the stronger result than Thm. 1: if there is no $3$-local quantity $\hat{Q}_{[3]}$ such that $\operatorname{\mathrm{len}}([\hat{Q}_{[3]},\hat{H}])\leq 2$, then there is no $k$-local quantity $\hat{Q}_{[k]}$ such that $\operatorname{\mathrm{len}}([\hat{Q}_{[k]},\hat{H}])\leq k-1$ for any $23\leq k\leq N/$ 2. Furthermore, the first isomorphisms (7) provide an insight into how to prove nonintegrability; first, we determine $\mathcal{B}_{\leq}^{(2)}$ and obtain all the elements of $\mathcal{B}_{\leq}^{(3)}$ by Eq. (7). Next, we search for the solution of $\operatorname{\mathrm{len}}([\hat{X}^{(3)}+\hat{X}^{(2)},\hat{H}])\leq 2$ for $\hat{X}^{(3)}\in\mathcal{B}_{\leq}^{(3)}$ and $\hat{X}^{(2)}\in\mathcal{B}^{(2)}\cap(\mathcal{B}^{(2)}_{\leq})^{\perp}$, where we remove the set of trivial solutions $\mathcal{B}^{(2)}_{\leq}\subset\mathcal{B}_{<}^{(3)}$. If there is no such nontrivial solution, we can conclude that the system is nonintegrable. This protocol not only unifies but also greatly simplifies the existing proofs of nonintegrability of systems satisfying (4[16, 19, 21, 22, 23, 24], where the absence of $k$-local conserved quantities is discussed separately for each $k$. A more specific procedure for translationally invariant systems, including how to determine $\mathcal{B}_{\leq}^{(2)}$, will be discussed in Sec. IV.
II.3 Example: nonintegrability of the XXZ chain with a non-uniform magnetic field
As an application of Thm. 1 and Thm. 2, we prove the absence of local conserved quantities in the spin-1/2 XXZ chain with a non-uniform magnetic field. The Hamiltonian of this system is represented as
@@ II.3 Example: nonintegrability of the XXZ chain with a non-uniform magnetic field
where $\hat{\sigma}^{\alpha}$’s represent the Pauli matrices. We assume $J_{x}=J_{y}\neq 0$ and $J_{z}\neq 0$. This is the same form as Eq. (3), where $\hat{H}^{(2)}_{i}$ is the XXZ interaction and $\hat{H}^{(1)}_{i}$ corresponds to a magnetic field. We refer to this magnetic field as uniform when $h_{i}$ takes the same value independently of $i$ ; otherwise, it is called non-uniform. If a magnetic field is uniform, this system is a Bethe solvable system [37], and local conserved quantities are explicitly constructed in Refs. [38, 39].
To consider local conserved quantities in this system, we first determine $\mathcal{B}_{\leq}^{(2)}$. In fact, this step has already been done in Refs. [16, 21, 22], and $\mathcal{B}_{\leq}^{(2)}=\mathbb{C}\hat{H}^{(2)}$ holds. By applying Eq. (7), we can determine $\mathcal{B}_{\leq}^{(k)}$; all of them are one-dimensional. In particular, $\mathcal{B}_{\leq}^{(3)}$ is generated by the following operator:
$\hat{Q}_{[3]}^{(3)}$ $\coloneqq\frac{\mathrm{i}}{2}\sum_{i=1}^{N}[\hat{H}^{(2)}_{i},\hat{H}^{(2)}_{i+1}]$
@@
Therefore, $c_{xxz,i}=c_{zyy,i-1}=0$ implies $h_{i+1}=h_{i}$, where we use $J_{x},J_{z}\neq 0$. This contradicts with non-uniformity of a magnetic field, and we have $\mathcal{B}_{<}^{(3)}\subset\mathcal{B}^{(2)}$. By applying Thm. 2, it turns out that $\mathcal{B}_{<}^{(k)}$ is also trivial for every $23\leq k\leq N/$ 2.
In summary, we have proven the following result.
@@
We comment on the connection between the above result and thermalization of this system. Many-body localization (MBL) is a phenomenon where strong disorder prevents a system from thermalizing, leading to a breakdown of the eigenstate thermalization hypothesis (ETH)  [40, 41, 42, 43]. Whether the XXZ model with a random magnetic field exhibits MBL in the thermodynamic limit is a subject of debate [44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. The MBL is attributed to the existence of an extensive number of quasi-local conserved quantities [44, 54], which are operators consisting of $k$-local operators ($0\leq k\leq N$) whose coefficients decay exponentially with $k$. Such a mechanism is sometimes referred to as the emergence of local integrability [42, 43], but note that this is different from the integrability we discuss, in the sense that there is an extensive number of (strictly) local conserved quantities. In fact, Prop. 3 tells us that the XXZ model with a random magnetic field is nonintegrable in our sense.
III Outline of Proof
III.1 String Diagram

### v3

@@ II.2 Main Result
Consider a Hamiltonian $\hat{H}$ with nearest-neighbor interactions and on-site potentials (see Eq. (3)) which satisfies the assumption (4). Then, we have the following linear isomorphisms between $\mathcal{B}_{\leq}^{(k)}$’s:
$\iota_{k}:\mathcal{B}_{\leq}^{(k-1)}\xrightarrow{\cong}\mathcal{B}_{\leq}^{(k)}$ (9)
for $23\leq k\leq N/$ 2, which acts on $\hat{X}^{(k)}=\sum_{i=1}^{N}\hat{X}^{(k-1)}_{i}$ in $\mathcal{B}_{\leq}^{(k-1)}$ as
$\hat{X}^{(k-1)}\mapsto\sum_{i=1}^{N}[\hat{X}^{(k-1)}_{i},\hat{H}^{(2)}_{i+k-2}]\in\mathcal{B}_{\leq}^{(k)}.$ (10)
Theorem 2 provides an insight into how to use Thm. 1 for proving nonintegrability; first, we determine $\mathcal{B}_{\leq}^{(2)}$ and check the assumption (6). Next, we search for the solution of $\operatorname{\mathrm{len}}([\iota_{2}(\hat{H}^{(2)})+\hat{X}^{(2)},\hat{H}])\leq 2$ for $\hat{X}^{(2)}\in\mathcal{B}^{(2)}$. If there is no such solution, we can conclude that the system is nonintegrable. This protocol not only unifies but also greatly simplifies the existing proofs of nonintegrability of systems satisfying (4) and (6[16, 21, 22, 23, 24], where the absence of $k$-local conserved quantities is discussed separately for each $k$. A more specific procedure for translationally invariant systems, including how to determine $\mathcal{B}_{\leq}^{(2)}$, will be discussed in Sec. V.
We note that the action of Eq. (10) is similar to that of the boost operator, which produces a tower of conserved quantities in Bethe solvable models [38, 39]. Even when assumption (6) fails, one can still prove the absence of conserved quantities generated by such boost from the Hamiltonian. By using the above isomorphisms, $\hat{H}^{(2)}\in\mathcal{B}_{\leq}^{(2)}$ is mapped to some operator
$\hat{Q}_{b}^{(k)}\coloneqq\iota_{k-1}\circ\dots\circ\iota_{2}(\hat{H}^{(2)})\in\mathcal{B}_{\leq}^{(k)}$ (11)
for $23\leq k\leq N/$ 2. These operators define a subspace of $\mathcal{B}_{<}^{(k)}$ as
$\mathcal{B}_{<,b}^{(k)}\coloneqq\mathcal{B}_{<}^{(k)}\cap(\mathbb{C}\hat{Q}_{b}^{(k)}\oplus\mathcal{B}^{(k-1)}),$ (12)
which represents the set of operator in $\mathcal{B}_{<}^{(k)}$ whose $k$-local component is proportional to $\hat{Q}_{b}^{(k)}$. Then, we have the following theorem.
Theorem 3.
Let us consider a Hamiltonian $\hat{H}$ of the same form as in Thm. 2. Then, we have the following linear isomorphisms between $\mathcal{B}_{<,b}^{(k)}$’s:
$\mathcal{B}_{<,b}^{(3)}\xrightarrow{\cong}\mathcal{B}_{<,b}^{(4)}\xrightarrow{\cong}\dots\xrightarrow{\cong}\mathcal{B}_{<,b}^{(\lfloor N/2\rfloor)},$ (13)
which map the subspace $\mathcal{B}_{\leq}^{(k-1)}\subset\mathcal{B}_{<,b}^{(k)}$ into $\mathcal{B}_{\leq}^{(k)}\subset\mathcal{B}_{<,b}^{(k+1)}$. In particular, if $\mathcal{B}_{<,b}^{(3)}=\mathcal{B}_{\leq}^{(2)}$ holds, then there is no $k$-local conserved quantity in $\mathcal{B}_{<,b}^{(k)}$ for $23\leq k\leq N/$ 2.
If the assumption (6) holds, we have $\mathcal{B}_{<,b}^{(k)}=\mathcal{B}_{<}^{(k)}$ and Thm. 3 yields the stronger result than Thm. 1: if there is no $3$-local quantity $\hat{Q}_{[3]}$ such that $\operatorname{\mathrm{len}}([\hat{Q}_{[3]},\hat{H}])\leq 2$, then there is no $k$-local quantity $\hat{Q}_{[k]}$ such that $\operatorname{\mathrm{len}}([\hat{Q}_{[k]},\hat{H}])\leq k-1$ for any $23\leq k\leq N/$ 2.
III Examples
As applications of Thm. 1 and Thm. 2, we prove the nonintegrability of several quantum spin systems.
III.1 The XXZ chain with a non-uniform magnetic field
We show the absence of local conserved quantities in the spin-1/2 XXZ chain with a non-uniform magnetic field. The Hamiltonian of this system is represented as
@@ III.1 The XXZ chain with a non-uniform magnetic field
where $\hat{\sigma}^{\alpha}$’s represent the Pauli matrices. We assume that $J_{x}=J_{y}\neq 0$ and $J_{z}\neq 0$. This model fits the form of Eq. (3), where $\hat{H}^{(2)}_{i}$ represents the XXZ interaction and $\hat{H}^{(1)}_{i}$ corresponds to the magnetic field. We refer to the magnetic field as uniform if $h_{i}$ is independent of $i$ ; otherwise, it is non-uniform. If a magnetic field is uniform, this system is a Bethe solvable system [40], and local conserved quantities are explicitly constructed in Refs. [41, 42].
To analyze the existence of local conserved quantities, we first determine $\mathcal{B}_{\leq}^{(2)}$. This step has been carried out in Refs. [16, 21, 22], and it was shown that $\mathcal{B}_{\leq}^{(2)}=\mathbb{C}\hat{H}^{(2)}$, which allows us to apply Thm. 1. By using Eq. (9), we can determine $\mathcal{B}_{\leq}^{(k)}$; all of them are one-dimensional. In particular, $\mathcal{B}_{\leq}^{(3)}$ is generated by the following operator:
$\hat{Q}_{[3]}^{(3)}$ $\coloneqq\frac{\mathrm{i}}{2}\iota_{2}(\hat{H}^{(2)})$
@@
Therefore, $c_{xxz,i}=c_{zyy,i-1}=0$ implies $h_{i+1}=h_{i}$, where we use $J_{x},J_{z}\neq 0$. This contradicts with non-uniformity of a magnetic field, and we have $\mathcal{B}_{<}^{(3)}\subset\mathcal{B}^{(2)}$. By applying Thm. 1, it turns out that there is no $k$-local conserved quantity for $23\leq k\leq N/$ 2.
In summary,we obtain the following result.
@@
We comment on the connection between the above result and thermalization of this system. Many-body localization (MBL) is a phenomenon where strong disorder prevents a system from thermalizing, leading to a breakdown of the eigenstate thermalization hypothesis (ETH)  [43, 44, 45, 46]. Whether the XXZ model with a random magnetic field exhibits MBL in the thermodynamic limit is a subject of debate [47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The MBL is attributed to the existence of an extensive number of quasi-local conserved quantities [47, 57], which are operators consisting of $k$-local operators ($0\leq k\leq N$) whose coefficients decay exponentially with $k$. Such a mechanism is sometimes referred to as the emergence of local integrability [45, 46], but note that this is different from the integrability we discuss, in the sense that there is an extensive number of (strictly) local conserved quantities. In fact, Prop. 4 tells us that the XXZ model with a random magnetic field is nonintegrable in our sense.
III.2 The $S=\frac{1}{2}$ XYZ model on the triangular lattice
To illustrate our main result in higher-dimensional systems, we consider the $S=\frac{1}{2}$ XYZ model on the triangular lattice. By embedding the triangular lattice into the square lattice (see Fig. 1), the Hamiltonian of the system takes the following form:
$\hat{H}=\sum_{\begin{subarray}{c}\mathbf{r}\in\Lambda_{N}\times\Lambda_{M}\\ \mathbf{v}=(1,0),(1,1),(0,1)\end{subarray}}\sum_{\alpha\in\{x,y,z\}}J_{\alpha}\hat{\sigma}_{\mathbf{r}}^{\alpha}\hat{\sigma}_{\mathbf{r}+\mathbf{v}}^{\alpha},$ (27)
where we assume that $M\geq 3$.
Figure 1: The triangular lattice can be regarded as the square lattice with additional edges. We define locality along the direction indicated by the horizontal lines.
We impose periodic boundary conditions (PBC) in one direction and either PBC or open boundary conditions (OBC) in the other:
$\hat{\sigma}_{(N+1,j)}^{\alpha}$ $=\hat{\sigma}_{(1,j)}^{\alpha},$
$\hat{\sigma}_{(i,M+1)}^{\alpha}$ =$=\begin{cases}\hat{\sigma}_{(i,1)}^{\alpha}\ (\mbox{PBC});\\ 0\ (\mbox{OBC}).\end{cases}$ (28)
We define locality by treating $\{i\}\times\Lambda_{M}$ as a single site in an effectively one-dimensional system. According to Eq. (3), the Hamiltonian is divided as
$\hat{H}_{i}^{(2)}$ $=\sum_{\begin{subarray}{c}j\in\Lambda_{M}\\ \mathbf{v}=(1,0),(1,1)\end{subarray}}\sum_{\alpha\in\{x,y,z\}}J_{\alpha}\hat{\sigma}_{(i,j)}^{\alpha}\hat{\sigma}_{(i,j)+\mathbf{v}}^{\alpha},$ (29)
$\hat{H}_{i}^{(1)}$ $=\sum_{j\in\Lambda_{M}}\sum_{\alpha\in\{x,y,z\}}J_{\alpha}\hat{\sigma}_{(i,j)}^{\alpha}\hat{\sigma}_{(i,j+1)}^{\alpha}.$ (30)
Based on the above definition, we have the following proposition.
Proposition 5.
The $S=\frac{1}{2}$ XYZ model on the triangular lattice (27) has no $k$-local conserved quantity for $23\leq k\leq N/$ 2.
Proof.
We prove Prop. 5 by using Thm. 1 and Thm. 2.
III.2.1 Injectivity of $\hat{H}_{i}^{(2)}$
Consider an operator $\hat{X}\in\mathcal{B}_{0}^{\{i\}\times\Lambda_{M}}$ such that $[\hat{X}\otimes I,\hat{H}_{i}^{(2)}]=0$. Our goal in this step is to show that $\hat{X}=0$. We expand $\hat{X}$ and $[\hat{X}\otimes I,\hat{H}_{i}^{(2)}]$ by the Pauli basis:
$\hat{X}=\sum_{\begin{subarray}{c}\bm{\alpha}\in\{0,x,y,z\}^{\Lambda_{M}}\\ \bm{\alpha}\neq\bm{0}\end{subarray}}q_{\bm{\alpha}}\bigotimes_{j\in\Lambda_{M}}\hat{\sigma}_{(i,j)}^{\alpha_{j}}\eqqcolon\sum_{\bm{\alpha}}q_{\bm{\alpha}}\hat{\bm{\sigma}}_{i}^{\bm{\alpha}},$ (31)
$[\hat{X}\otimes I,\hat{H}_{i}^{(2)}]=\sum_{\begin{subarray}{c}\bm{\alpha}\\ \beta\in\{x,y,z\}\\ j^{\prime}\in\Lambda_{M}\end{subarray}}r_{\bm{\alpha},\beta;j^{\prime}}\hat{\bm{\sigma}}_{i}^{\bm{\alpha}}\hat{\sigma}_{(i+1,j^{\prime})}^{\beta},$ (32)
where $\hat{\sigma}_{0}\coloneqq I$ is the identity operator. When we consider the OBC, we take $\alpha_{0}=\alpha_{M+1}=0$.
First, we show $q_{\bm{\alpha}}=0$ if $\alpha_{j}=0$ for some $j\in\Lambda_{M}$. Since $\bm{\alpha}\neq\bm{0}$, we can appropriately redefine $j$ so that either $\alpha_{j+1}\neq 0$ or $\alpha_{j-1}\neq 0$ holds. If $\alpha_{j+1}=x$, then we have
$r_{\bm{\alpha}^{\prime},y;j+1}=2\mathrm{i}J_{y}q_{\bm{\alpha}},$ (33)
where
$\alpha^{\prime}_{j^{\prime}}=\begin{cases}\alpha_{j^{\prime}}&(j^{\prime}\neq j+1)\\ z&(j^{\prime}=j+1)\end{cases}.$ (34)
Therefore, $[\hat{X}\otimes I,\hat{H}_{i}^{(2)}]$ implies $q_{\bm{\alpha}}=0$. The remaining cases follow similarly.
Next, we consider $\bm{\alpha}\in\{x,y,z\}^{\Lambda_{M}}$ such that there exist $j,j^{\prime}\in\Lambda_{M}$ satisfying $\alpha_{j}\neq\alpha_{j^{\prime}}$. In this case, we can take some $1\leq j<M$ satisfying $\alpha_{j}\neq\alpha_{j+1}$. If $(\alpha_{j},\alpha_{j+1})=(y,x)$, we have $r_{\bm{\alpha}^{\prime},y;j+1}=2iJ_{y}q_{\bm{\alpha}}$ for $\bm{\alpha}^{\prime}$ in Eq. (34). The same applies to other cases, and we have shown $q_{\bm{\alpha}}=0$ if $\alpha_{j}\neq\alpha_{j^{\prime}}$ for some $(j,j^{\prime})$.
Finally, we examine the case where $\alpha_{j}=\alpha_{1}\in\{x,y,z\}$ for all $j\in\Lambda_{M}$. If $\alpha_{1}=x$, we have
$r_{\bm{\alpha}^{\prime},y;2}=2\mathrm{i}J_{y}(q_{\bm{\alpha}}-q_{\tilde{\bm{\alpha}}}),$ (35)
where $\bm{\alpha}^{\prime}$ is defined in Eq. (34) and
$\tilde{\alpha}_{j}=\begin{cases}z&(j=1,2)\\ x&(j\neq 1,2)\end{cases}.$ (36)
Therefore, $[\hat{X}\otimes I,\hat{H}_{i}^{(2)}]$ implies $q_{\bm{\alpha}}=q_{\tilde{\bm{\alpha}}}$. On the other hand, we have shown $q_{\tilde{\bm{\alpha}}}=0$ in the above. Thus, $q_{\bm{\alpha}}=0$. The same argument can be made for $\alpha_{1}=y,z$, and it follows that $q_{\bm{\alpha}}=0$. From the above, it follows that $\hat{X}=0$, and the first line of Eq. (4) is proven. The proof for the second line can be done in a similar manner.
III.2.2 Confirmation of Assumption (6)
In this step, we confirm the assumption (6), or more specifically, we show that $\hat{X}\propto\hat{H}^{(2)}$ for $\hat{X}\in\mathcal{B}_{\leq}^{(2)}$. First, we note that $\hat{X}\in\mathcal{B}_{\leq}^{(2)}$ is equivalent to
$[\hat{X}_{i}\otimes I,I\otimes\hat{H}^{(2)}_{i+1}]=[\hat{H}^{(2)}_{i}\otimes I,I\otimes\hat{X}_{i+1}]$ (37)
for all $i\in\Lambda_{N}$. From this equation and the injectivity, we know 333 See the discussion at the last paragraph of Sec. V. In this case, $\mathcal{A}$ can be taken as the subspace consisting of one-body operators. that $\hat{X}_{i}$ consists of two-body operators:
$\hat{X}_{i}=\sum_{\begin{subarray}{c}\alpha,\beta\in\{x,y,z\}\\ j,k\in\Lambda_{M}\end{subarray}}q_{\alpha\beta;jk,i}\hat{\sigma}^{\alpha}_{(i,j)}\hat{\sigma}^{\beta}_{(i+1,k)}.$ (38)
Furthermore, it is clear that $q_{\alpha\beta;jk,i}=0$ unless $k\in\{j,j+1\}$.
Next, we show that $q_{\alpha\beta;jk,i}=q\delta_{\alpha\beta}J_{\alpha}$ for some $q\in\mathbb{C}$. By comparing the coefficient of $\hat{\sigma}^{x}_{(i,j)}\hat{\sigma}^{x}_{(i+1,k)}\hat{\sigma}^{z}_{(i+2,k)}$ on both sides of Eq. (37), we have $02iJ_{z}q_{xy;jk,i}=$ 0. Similarly, we can show that $q_{\alpha\beta;jk,i}=0$ for $\alpha\neq\beta$. To consider the case of $\alpha=\beta$, we compare the coefficient of $\hat{\sigma}^{x}_{(i,j)}\hat{\sigma}^{z}_{(i+1,k)}\hat{\sigma}^{y}_{(i+2,l)}$ on both sides of Eq. (37). This yields relations between $q_{\alpha\alpha;jk,i}$’s as
$J_{y}q_{xx;jk,i}=J_{x}q_{yy;kl,i+1}$ (39)
for all $i\in\Lambda_{N}$, $k\in\{j,j+1\}\cap\Lambda_{M}$ and $l\in\{k,k+1\}\cap\Lambda_{M}$. Similarly, we have
$q_{\alpha\alpha;jk,i}/J_{\alpha}=q_{\beta\beta;kl,i+1}/J_{\beta},$ (40)
which is equivalent to $q_{\alpha\alpha;jk,i}=\frac{q_{xx;11,1}}{J_{x}}J_{\alpha}$ for all $i\in\Lambda_{N}$ and $k\in\{j,j+1\}\cap\Lambda_{M}$. Hence, we have shown that $\hat{X}=\frac{q_{xx;11,1}}{J_{x}}\hat{H}^{(2)}$. Therefore, the Hamiltonian (27) satisfies the assumption (6).
III.2.3 Absence of $3$-local quantity $\hat{Q}_{[3]}$ satisfying $\operatorname{\mathrm{len}}([\hat{Q}_{[3]},\hat{H}])\leq 2$
Up to this point, we have confirmed that the Hamiltonian (27) satisfies the assumption (4) and (6). Thus, Thm. 1 can be applied, and it suffices to show that there is no $3$-local quantity $\hat{Q}$ such that $\operatorname{\mathrm{len}}([\hat{Q},\hat{H}])\leq 2$ holds. By the fact (6) and Thm. 2, a $3$-local quantity $\hat{Q}_{[3]}$ satisfying $\operatorname{\mathrm{len}}([\hat{Q}_{[3]},\hat{H}])\leq 2$ can be written as
$\hat{Q}_{[3]}=q\sum_{i=1}^{N}[\hat{H}_{i}^{(2)},\hat{H}_{i+1}^{(2)}]+\hat{Q}_{[3]}^{(2)}$ (41)
with some $q\in\mathbb{C}$ and 2-local quantity $\hat{Q}_{[3]}^{(2)}$. Our goal is to show that $q=0$.
We expand this operator $\hat{Q}_{[3]}^{(2)}$ by the Pauli basis:
$\hat{Q}_{[3],i}^{(2)}=\sum_{\bm{\alpha},\bm{\beta}\in\{0,x,y,z\}^{\Lambda_{M}}}q^{(2)}_{\bm{\alpha}\bm{\beta},i}\hat{\bm{\sigma}}^{\bm{\alpha}}_{i}\hat{\bm{\sigma}}^{\bm{\beta}}_{i+1}.$ (42)
We focus on the following $(\bm{\alpha},\bm{\beta})$:
$\alpha_{j}$ =$=\begin{cases}z&(j=1)\\ x&(j=2)\\ 0&(j\neq 1,2)\end{cases},$ (43)
$\beta_{j}$ =$=\begin{cases}y&(j=1)\\ 0&(j\neq 1)\end{cases}.$ (44)
The coefficient of $\hat{\sigma}_{(1,1)}^{x}\hat{\sigma}_{(2,1)}^{y}\hat{\sigma}_{(2,2)}^{x}\hat{\sigma}_{(3,1)}^{y}$ in $[\hat{Q}_{[3]},\hat{H}]$ comes from the following commutators:
$[q^{(2)}_{\bm{\alpha}\bm{\beta},2}\hat{\sigma}_{(2,1)}^{z}\hat{\sigma}_{(2,2)}^{x}\hat{\sigma}_{(3,1)}^{y},J_{x}\hat{\sigma}_{(1,1)}^{x}\hat{\sigma}_{(2,1)}^{x}],$ (45)
$[2\mathrm{i}qJ_{x}J_{y}\hat{\sigma}_{(1,1)}^{x}\hat{\sigma}_{(2,1)}^{z}\hat{\sigma}_{(3,1)}^{y},J_{x}\hat{\sigma}_{(2,1)}^{x}\hat{\sigma}_{(2,2)}^{x}].$ (46)
Since the coefficient of $\hat{\sigma}_{(1,1)}^{x}\hat{\sigma}_{(2,1)}^{y}\hat{\sigma}_{(2,2)}^{x}\hat{\sigma}_{(3,1)}^{y}$ is zero by the assumption $\operatorname{\mathrm{len}}([\hat{Q}_{[3]},\hat{H}])\leq 2$, the following equation holds:
$q^{(2)}_{\bm{\alpha}\bm{\beta},2}+2\mathrm{i}qJ_{x}J_{y}=0.$ (47)
Similarly, by focusing on the coefficient of $\sigma_{(2,1)}^{z}\sigma_{(2,2)}^{x}\sigma_{(3,1)}^{z}\sigma_{(4,1)}^{x}$, we have
$q^{(2)}_{\bm{\alpha}\bm{\beta},2}-2\mathrm{i}qJ_{x}J_{y}=0.$ (48)
By combining these equations, we have $q=0$. Therefore, there is no $3$-local quantity $\hat{Q}_{[3]}$ satisfying $\operatorname{\mathrm{len}}([\hat{Q}_{[3]},\hat{H}])\leq 2$ and this system has no $k$-local conserved quantity for $23\leq k\leq N/$ 2. ∎
IV Outline of Proof
IV.1 String Diagram